The average number of days school is closed due to snow during the winter in a certain city in the eastern part of United States is 4.What is the probability that the school in this city will close for 2 or more days during winter
Answer by stanbon(75887)

The average number of days school is closed due to snow during the winter in a certain city in the eastern part of United States is 4.What is the probability that the school in this city will close for 2 or more days during winter
-------------------------------------
You cannot answer this unless you know the probability of snow on a day
in the winter.
It's a binomial problem with np = 4.
If you assume the "winter" means 365/4 = 91 days then p = 4/91.
In that case P(x >= 2) = 1 - binomcdf(91,(4/91),1) = 0.9133
-------------------------------------------------------------------
Cheers,
Stan H.